OpenStudy (anonymous):

Solve the system using the addition method. 3x + 2y = 10 6x + 4y = 14 What is the solution to the system?

OpenStudy (anonymous):

x = (10-2y)/3 6(10-2y)/3 + 4y = 14 20 - 4y + 4y = 14 the given is incorrect.

OpenStudy (anonymous):

Ahi! The above means that 20=14... (multiply the first equation by 2, and compare to the second) i.e. there are no solutions

OpenStudy (anonymous):

wait what? im confused

myininaya (myininaya):

the lines are parallel they have no intersection

OpenStudy (anonymous):

ooh